Prev Next

SW&IT Meeting Minutes 2022-11-28

Attendees: Brandon, David, Glenn, Matthieu, Paul, Stephen, Xianyu, Yang

Outlook

Labor Reporting

Report auto-generated every Monday at 6:30am, Do not edit this table, Update your entries in smartsheets.

20221114
LR20221114 - Bec, Matthieu 64.0
LR20221114 - Mechtley, Brandon 56.0
LR20221114 - Eychaner, Glenn 60.0
LR20221114 - Edmeades, David 64.0
LR20221114 - Grenz, Paul 0.0
LR20221114 - Hooper, Stephen 24.0
LR20221114 - Zhang, Yang 40.0
LR20221114 - Zhang, Xianyu 41.0

News from management

  • Please submit your Career Conversation forms in UAccess for 2022

Previous Action Items

Owner Action Due Date Completion Date

Open topics

  • IT 8774 (Non-sidereal tracking FITS headers) - need agreement on FITS header keyword
  • Re: disks for AO computers - aka, ADSEC machines
  • LBTI/AHS - it's broken, use the simpler native CA approach
  • TCS snafus - dumbo dev releases, pmac client*s*, OSS rpc log floods, netconfig restart, etc - QA ???
  • tcs-test, good to recycle as spare for mountain?
  • wfs spare status for SOUL handover - centos6, mountain relocation?
  • LBTI wfs DNS update
  • obs2 /var/log/messages is 2G and indicates some log rotation is not happening. And flatpack ?
  • Network problems at the telescope (maybe) AdSec are in failure

Team reports

MB

  • Last week
    • OC, Management
    • playing with tcs VM
    • AHS / slack release
    • trying to overcome turkey OD (it has a name that I cannot remember)
  • This week
    • tcs sim VM
    • AOS patch PR #209
    • obs station iteration, 5x in 507
    • dms lttb
    • SOUL team visit
    • look at lib-telemetry and h5 patch (needed or not)

YZ

  • Will try to schedule some TMS test if possible.
  • Update TMS code to use the "epics" module.
  • Release new RB_Science program.
  • Release new OSCO program and rebuild the RPM.
  • Updating archive utility.
  • Start on some new utilities (discussed with Olga), adding differential tracking rate info to LBC/MODS images. Currently planning to use the same approach as the MODS WCS program (e.g., build special database to capture TCS information, then extract those information and add them to images in post-processing steps on the archive machine).
  • Also need to work on adding TMS information to LBC images.
  • Maybe work on some TCS stuff as well.

DE

  • Last week
    • Arranged a ride for the spare WFS server, Josh is taking it up with him tomorrow
    • Finally got the failed 18TB disk shipped out
    • Archive-in newdata hiccup; reboot fixed it.
      • Will research error and see if I can pinpoint the cause
    • Upgrade verification tests on XZ's ao-test machine.
  • This week
    • Guido Agapito VPN? I'll ask him to fill the Google Sheet for his details
    • Argos laptop, whenever that makes it to Tucson
    • Sundries as people get back into the swing

XZ

  • Last week
    • ASM housekeeper reads from MOXA wiki and preparaton
    • ASM support
    • LN slopes and DX AdSec discussion and test plan
    • spase disks for AdSec with David
    • AdSec DSP watchdog expire
  • This week
    • HK test
    • IT8771 (RR modes inaccurately reported in AOS) test
    • work with SOUL team
    • LN SX GWFS close loop test
    • ASM/AO support

GCE

  • Last week:
    • IT 8778 (RTD opening files): J.Power noted that the RTD can't open files in /lbt/data/new or /repository directly. This is by design, since Aladin wants write access to those directories (and any directory it reads files from). The UNIX file and mount permissions probably won't let Aladin overwrite data, but I deliberately gave Aladin the absolute minimum SecurityManager permissions to run the RTD.
    • Discovered that the Observer GUI and every other GUI in the LUCI software make the same mistakes regarding concurrency, threading, and modifying GUI elements outside the Event thread as the RTD did, though without consequences quite as drastic (though this may be why the Observer GUI occasionally freezes). Decided that the solutions I used in the RTD would apply just as well elsewhere, and refactored the threading and concurrency code so that it could be used without a dependency on Aladin.
    • Out of office all Tuesday afternoon.
  • This week:
    • Install the GEIRS and LUCI software onto the new LUCI computer.
    • More Observer GUI recoding and testing.
Topic revision: r16 - 28 Nov 2022, DavidEdmeades
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback